# Safe Reinforcement Learning on the Constraint Manifold: Theory and Applications

### v1

@@ Abstract
Integrating learning-based techniques, especially reinforcement learning, into robotics is promising for solving complex problems in unstructured environments. However, most existing approaches are trained in well-tuned simulators and subsequently deployed on real robots without online fine-tuning. In this setting, the simulation’s realism seriously impacts the deployment’s success rate. Instead, learning with real-world interaction data offers a promising alternative: not only eliminates the need for a fine-tuned simulator but also applies to a broader range of tasks where accurate modeling is unfeasible. One major problem for on-robot reinforcement learning is ensuring safety, as uncontrolled exploration can cause catastrophic damage to the robot or the environment. Indeed, safety specifications, often represented as constraints, can be complex and non-linear, making safety challenging to guarantee in learning systems. In this paper, we show how we can impose complex safety constraints on learning-based robotics systems in a principled manner, both from theoretical and practical points of view. Our approach is based on the concept of the Constraint Manifold, representing the set of safe robot configurations. Exploiting differential geometry techniques, i.e., the tangent space, we can construct a safe action space, allowing learning agents to sample arbitrary actions while ensuring safety. We demonstrate the method’s effectiveness in a real-world Robot Air Hockey task, showing that our method can handle high-dimensional tasks with complex constraints. Videos of the real robot experiments are available on the project website [^1] https://puzeliu.github.io/TRO-ATACOM.
@@ I Introduction
Figure 2: The robot air hockey task. The objective is to strike the puck to the opponent’s goal. The vector field shows the velocity of the end-effector at different locations when a positive unit action is applied in the first two dimensions of the safety action space using ATACOM. The blue (resp. red) arrow corresponds to a unit action applied in the first (resp. second) dimension.missingmissing
@@
Theoretically, this online adaptation is feasible using the Reinforcement Learning (RL) framework, which describes an optimization technique for the policy under a performance metric defined as a reward function. However, many key problems are preventing the usage of online RL techniques in the real world. On the one hand, this technique often requires an unreasonable large amount of data to learn the desired policy and to adapt to environmental variations. On the other hand, the learning process can arbitrarily change the policy and require explorative actions, possibly causing the robotic system to take dangerous actions that may harm people or cause damage to the environment or to the robot itself. While the effort of the RL, in general, and the Robot Learning community, in particular, has primarily focused on improving the efficiency of the learning algorithm in challenging environments, the importance of acting under safety constraints keeps increasing. Complying with safety constraints is fundamental for allowing robots to operate in the real world, but, unfortunately, imposing safety constraints on a learning system is not as straightforward as in a classical robotic system.
@@
Among many techniques to enforce safety constraints on learning agents, the most popular one is to exploit the Safe Reinforcement Learning (SafeRL) framework. This framework is based on the Constrained Markov Decision Processes (CMDP) formalism [6], and the algorithm objective is to ensure the learned policy fulfills the safety constraints while maximizing the reward function. By construction, these methods do not ensure the fulfillment of safety constraints during the learning process; they only converge to a final safe policy. Therefore, these algorithms are unsuitable for online adaptation of the learning agents in the real world, where any safety constraint violation should be avoided during the training process. To tackle this problem, researchers developed Safe Exploration approaches to ensure safety during the whole learning process [7, 8]. Unfortunately, performing safe exploration without relying on domain knowledge is impossible. Pioneer researchers have exploited different types of prior domain knowledge to obtain a safe exploration policy, such as pre-defined safety constraints, previous data demonstrating the safe and unsafe regions, dynamics models, or pre-defined safety fallback controllers. Choosing how much and what domain knowledge is necessary is challenging, as it is a trade-off between safety, performance, and design time.
@@
In this paper, we extend and analyze in depth the method, ATACOM, which was proposed in [9] and extended in [10]. ATACOM is a simple but effective approach for achieving safe exploration while using minimal knowledge of the robotic platform. ATACOM is based on the concept of Constraint Manifold [11, 12], which converts the constraint optimization to an optimization problem on the manifold. We build a safe action space on the tangent space of the constraint manifold, leveraging the known differential dynamics of the robotic system and tools from differential geometry. Using this approach, any action sampled from the safe action space will keep the system on the constraint manifold, ensuring safety. While the core methodology was already presented in previous works, in this paper, we extend the ATACOM framework with four novel contributions:
@@
    We provide a novel formulation and a theoretical analysis of the ATACOM method. Using LaSalle’s principle and under some mild conditions, we prove that the proposed method guarantees the system’s safety.
@@
  4. 4.
    We show in experiments the scalability and robustness of ATACOM to model mismatch in a high-dimensional complex environment, the Robot Air Hockey (7-dimensional action with 21 constraints) [13], and finally demonstrate the effectiveness of the method for online adaptation in the real world.
@@ II Related Work
In the last decades, CMDP [14, 6] has attracted a lot of interest from RL researchers in solving constrained control problems. Under this framework, several different forms of constraints have been studied. One important form of constraint is the expected cumulative cost. The objective is formulated as maximizing the expected return while maintaining the expected cost below a threshold [15, 16, 17, 18, 19, 20, 21, 22, 23]. This type of constraint has been extended to different variants, such as the risk-sensitive constraint [24, 25, 26, 27] and the probabilistic constraint [28, 29, 30]. Different types of constrained optimization techniques are applied in the policy update process, such as the trust-region method [15, 26], the interior point method [18], and the Lagrangian relaxation method [14, 17, 19, 20, 24, 25, 27]. Furthermore, the Lyapunov function is also used to derive a policy improvement procedure [16, 31, 32]. Although these methods focus on obtaining a safe policy at the end of the training, they often violate the constraints during the learning process. Such methods are unsuitable for learning directly on real robots or for online fine-tuning, as they usually rely on a value function or environment model, and model mismatches can destroy the safety value function.
Another type of approach focuses on adhering to state-dependent constraints at every time step, often referred to as the safe exploration method. The safe exploration problem consists of two significant challenges: (1) construction of safe constraints [33] and (2) obtaining safe action. The constraint should account for the safety of not only the current state but also the future trajectories. Control barrier functions [34, 35, 36, 37] and reachability analysis [38, 39, 40, 41] utilize prior knowledge of the system dynamics or learned dynamics model to construct the safety constraint. Chance constraint approaches analyze the safety problem from a probabilistic perspective, i.e., the probability of encountering an unsafe state is below a threshold [42, 43, 44]. Temporal logic constraints [45, 46] provide a symbolic expression for the notion of safety. In this paper, we do not restrict the type of constraints as long as they are differentiable. Instead, we focus on the challenge of obtaining safe actions.
@@
The most common choice for determining the safe action at each time step is to add a safety layer on top of the policy, which will correct the unsafe action to a safe one [47, 48, 49]. Nevertheless, determining an action that satisfies the safety constraint or corrects the constraint violation is almost impossible without incorporating prior knowledge. Exploiting the dynamics model to find the safe action has been applied to safe-critical control and learning tasks [34, 50]. Defining a task-specific backup policy enables safe exploration in low-dimensional tasks, such as navigation and pendulum [7, 51, 52]. Learning dynamic models using offline datasets and deriving safe actions using the learned models are applicable to tasks where dynamic models are unavailable or inaccurate [53, 54, 37, 49]. Our approach falls into the safe exploration category, aiming to find a safe action at each time step. Our method utilizes the knowledge of the dynamics model to construct a safe action space. Unlike existing approaches, our method does not require an initial safe policy, a backup policy, or solving a constrained optimization problem. ATACOM constructs a safe action space where all sampled actions are guaranteed to be safe. Therefore, our method is not explicitly restricted to any learning algorithm.
@@
Safe learning for robotics, as the core intersection between the SafeRL and robotics [55], has been raised for various types of application scenarios, varying from manipulation [56, 57] and navigation [58, 59], to locomotion [60] and Human-Robot Interaction (HRI) [61, 62, 63, 64]. Previous work of ATACOM has demonstrated the effectiveness of the method in manipulation, navigation, and HRI tasks [9, 10]. In this paper, we focus on a better illustration of individual components with simple low-dimensional environments and finally verify the method in a high-dimensional real-world robotic task.
III Preliminaries
III-A Differential Geometry
@@
$\mathrm{T}_{x}\mathcal{M}$ $=\mathrm{ker}\;\mathrm{D}f({\bm{x}})$
=$=\left\{{\bm{v}}\in\mathcal{E}:\langle\mathrm{grad}f_{i}({\bm{x}}),{\bm{v}}\rangle=0,i\in\{1,\dots,k\}\right\}$
The dimension of the tangent space is the same as the given manifold $\mathrm{dim}{\mathcal{M}}=d-k$.
III-B LaSalle’s Invariance Principle
We present the LaSalle’s Invariance Principle [66], which is used later in the discussion to prove the safety. Consider the autonomous system
@@ III-B LaSalle’s Invariance Principle
III-C Problem Statement
In SafeRL, we model the environment as a CMDP. A CMDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},P,\gamma,R,\mu_{0},\mathcal{K}\rangle$ with the state space $\mathcal{S}$, the action space $\mathcal{A}$, the state transition probability kernel $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$, the discount factor $\gamma\in(0,1]$, the reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, the initial state distribution $\mu_{0}:\mathcal{S}\rightarrow\mathbb{R}^{+}$, and the set of constraint functions $\mathcal{K}:=\{k_{i}:\mathcal{S}\rightarrow\mathbb{R},i\in\{1,\dots,K\}\}$. In this paper, we approach the safety problem in RL from the point of view of Safe Exploration. In this setting, the objective is to prevent constraint violations throughout the whole learning process. This setting corresponds to the following constrained optimization problem:
@@ III-C Problem Statement
with the trajectory $\tau=[{\bm{s}}_{0},{\bm{a}}_{0},\dots,{\bm{s}}_{T}]$, the state ${\bm{s}}_{t}\in\mathcal{S}\subset\mathbb{R}^{S}$, and the action ${\bm{a}}_{t}\in\mathcal{A}\subset\mathbb{R}^{A}$ sampled from the policy $\pi$. In this formulation, the inequality constraints $k_{i}:\mathcal{S}\rightarrow\mathbb{R}$ specify the safety requirements at each time step among all trajectories.
Modern RL algorithms solve this optimization by evaluating a sequence of policies $\{\pi_{0},\dots\pi_{j}\}$ in the environment, where the learning algorithm produces a new policy $\pi_{j+1}$ using $\mathcal{D}_{j}$, i.e. the dataset of environment interactions using the previous policies $\{\pi_{0},\dots,\pi_{j}\}$. In this scenario, Safe Exploration requires that every state ${\bm{s}}$ from every trajectory $\tau$ sampled from each policy $\pi_{j}$ is safe, i.e., $k_{i}({\bm{s}})\leq 0,\forall i$. This requires that the optimization algorithm selects each policy $\pi_{j}$ from $\Pi_{\text{safe}}$, i.e. the space of policies respecting the safety constraints. Most safe exploration approaches design a space $\tilde{\Pi}\subseteq\Pi_{\text{safe}}$ and solve the optimization problem in (3) generating $\pi_{j}\in\tilde{\Pi}$.
In the rest of this work, we focus on the control affine system that is local Lipschitz continuous. The system is described as follows:
@@ IV Methods
In this section, we introduce the technical details of our proposed approach ATACOM and discuss under which conditions the designed controller is safe. Briefly, our approach tries to construct a safe action space tangent to the Constraint Manifold. Combining the basis of the tangent space with a coordinate determined by the controller, that can be hand-crafted or an arbitrary RL method, we can guarantee the safety of the overall control action. The conceptual illustration of the approach is shown in Fig. 3. Our approach is particularly suitable for RL as the actions are sampled randomly from a safe space, and the policy is safe by construction.
As it is fundamental for our approach, we will first rigorously define the Constraint Manifold and the Tangent Space at each point on the manifold. Then, we introduce an augmented state dynamic model that incorporates the dynamics of the slack variable. Next, we derive a safe controller that moves on the constraint manifold equipped with augmented dynamics. We then theoretically analyze the conditions of the controller to guarantee safety. At the end of this section, we extend our safety proof from a Lipschitz continuous system to switched systems suitable for RL setup.
@@ IV-A Definition of the Constraint Manifold
The constraint function $k:\mathcal{S}\rightarrow\mathbb{R}^{K}$ is of class $C^{1}$.
@@
Consider the inequality constraint $k(s)\coloneqq s\leq 0$. The constraint manifold =$\mathcal{M}=\{(s,\mu)\in\mathbb{R}\times[0,\infty):c(s,\mu)\coloneqq s+\mu=0\}$ is a 1-dimensional submanifold embedded in $\mathbb{R}^{2}$. The constraint Jacobian is $J_{c}(s,\mu)=[1\quad 1]$. We can choose the basis of the tangent space as $B(s,\mu)=\frac{1}{\sqrt{2}}[1,-1]^{\intercal}$.

### v2

@@ Abstract
Integrating learning-based techniques, especially reinforcement learning, into robotics is promising for solving complex problems in unstructured environments. However, most existing approaches are trained in well-tuned simulators and subsequently deployed on real robots without online fine-tuning. In this setting, extensive engineering is required to mitigate the sim-to-real gap, which can be challenging for complex systems. Instead, learning with real-world interaction data offers a promising alternative: it not only eliminates the need for a fine-tuned simulator but also applies to a broader range of tasks where accurate modeling is unfeasible. One major problem for on-robot reinforcement learning is ensuring safety, as uncontrolled exploration can cause catastrophic damage to the robot or the environment. Indeed, safety specifications, often represented as constraints, can be complex and non-linear, making safety challenging to guarantee in learning systems. In this paper, we show how we can impose complex safety constraints on learning-based robotics systems in a principled manner, both from theoretical and practical points of view. Our approach is based on the concept of the Constraint Manifold, representing the set of safe robot configurations. Exploiting differential geometry techniques, i.e., the tangent space, we can construct a safe action space, allowing learning agents to sample arbitrary actions while ensuring safety. We demonstrate the method’s effectiveness in a real-world Robot Air Hockey task, showing that our method can handle high-dimensional tasks with complex constraints. Videos of the real robot experiments are available publicly [^1] https://puzeliu.github.io/TRO-ATACOM.
@@ I Introduction
Theoretically, this online adaptation is feasible using the Reinforcement Learning (RL) framework, which describes an optimization technique for the policy under a reward function. However, many key problems are preventing the usage of online RL techniques in the real world. On the one hand, RL method often requires an unreasonable large amount of data to learn the desired policy and to adapt to environmental variations. On the other hand, the learning process can arbitrarily change the policy and require explorative actions, possibly causing the robotic system to take dangerous actions that may harm people or cause damage to the environment or to the robot itself. While the effort of the RL, in general, and the Robot Learning community, in particular, has primarily focused on improving the efficiency of the learning algorithm, the importance of acting under safety constraints keeps increasing. Complying with safety constraints is fundamental for allowing robots to operate in the real world, but, unfortunately, imposing safety constraints on a learning system is not as straightforward as in a classical robotic system.
@@
Among many techniques to enforce safety constraints on learning agents, the most popular one is to exploit the Safe Reinforcement Learning (SafeRL) framework. This framework is based on the Constrained Markov Decision Processes (CMDP) formalism [6], and the objective is to ensure the learned policy fulfills the safety constraints while maximizing the reward function. By construction, these methods do not ensure the fulfillment of safety constraints during the learning process; they only converge to a final safe policy. Therefore, these algorithms are unsuitable for online adaptation of the learning agents in the real world, where any safety constraint violation should be avoided during the training process. To address this challenge, researchers have developed Safe Exploration approaches that ensure safety throughout the entire learning process [7, 8] by integrating prior knowledge, such as robot dynamics [9], pre-defined constraints [10], or demonstrations [11], to derive effective safety constraints. Once these constraints are established, safety is typically enforced using a safety filter. This filter can involve solving a constrained optimization problem, such as a quadratic program (QP), or substituting unsafe actions with pre-defined backup actions. Alternatively, ATACOM provides a novel approach that constructs a safe action space, eliminating the need for solving a constrained optimization iteratively or designing a backup policy.
@@
In this paper, we extend and analyze in depth the method, ATACOM, which was proposed in [12] and extended in [13]. ATACOM is a simple but effective approach for achieving safe exploration exploiting the knowledge of the robot dynamics and constraints. ATACOM is based on the concept of Constraint Manifold [14, 15], which converts the constraint optimization to an optimization problem on the manifold. We build a safe action space on the tangent space of the constraint manifold, leveraging the known dynamics of the robotic system and tools from differential geometry. Using this approach, any action sampled from the safe action space will keep the system on the constraint manifold, ensuring safety. In this paper, we extend the ATACOM method with following contributions:
@@
    We provide a new formulation of the method by constructing a dynamic model for the slack variable. Despite the mathematical equivalence, the new formulation provides an interpretation from the dynamical system perspective, allowing us to conduct a theoretical analysis using LaSalle’s principle and input-to-state stability.
@@
    We analyze the ATACOM controller with the existence of disturbances, showing that ATACOM controller ensures bounded constraint violations with bounded disturbances.
  3. 3.
    We present a novel method for constructing a smoothly varying tangent basis, effectively addressing the issue of discontinuous bases.
  4. 4.
@@
  6. 6.
    For the first time, we demonstrate that ATACOM can perform online fine-tuning of the policy in a contact-rich, dynamic task (Robot Air Hockey) safely, whereas former works on ATACOM only deployed pre-trained policies on real robots.
@@ II Related Work
In the last decades, CMDP [16, 6] have garnered significant attention from RL researchers for addressing constrained control problems. Various forms of constraints have been explored within this framework. One prominent constraint is the expected cumulative cost, where the objective is to maximize the expected return while ensuring that the expected cost remains below a specified threshold [17, 18, 19, 20, 21, 22, 23, 24, 25]. This constraint has been adapted into different variants, including risk-sensitive constraint [26, 27, 28, 29] and the probabilistic constraint [30, 31, 32]. Several optimization techniques have been applied to update policies under constraints, such as the trust-region method [17, 28], the interior point method [20], Lagrangian relaxation [16, 19, 21, 22, 26, 27, 29], and Lyapunov function-based methods [18, 33, 34]. While these approaches aim to derive a safe policy by the end of training, they often violate constraints during the learning process, making them unsuitable for direct application on real robots or for online fine-tuning tasks.
To address these shortcomings, Safe Exploration (SafeExp) focuses on satisfying state-dependent constraints at every time step, and it presents two significant challenges: (1) constructing safe constraints [35] and (2) obtaining safe action. When constructing constraints, they should account for the safety of both the current state and future trajectories. To address the first challenge, methods such as control barrier functions [10, 36, 37, 38] and reachability analysis [39, 40, 41, 9, 42] utilize prior knowledge of the system dynamics or learned dynamics model to construct the safety constraint. To alleviate the engineering effort of manually designing constraints, learning-based techniques have been applied to learn a safety constraint [43, 44]. In this paper, we assume the presence of state-dependent constraints and focus on the challenge of obtaining safe actions.
@@
To obtain a safe actions, the most common choice is to add a safety layer on top of the policy, which will correct the unsafe action to a safe one [45, 46, 47]. Nevertheless, determining an action that satisfies the safety constraint or corrects the constraint violation is almost impossible without incorporating prior knowledge. Exploiting the dynamics model to find the safe action has been applied to safe-critical control and learning tasks [10, 48]. Alternatively, defining a task-specific backup policy enables safe exploration in low-dimensional tasks, such as navigation and pendulum [7, 49, 50]. Learning dynamic models using offline datasets and deriving safe actions using the learned models are applicable to tasks where dynamic models are unavailable or inaccurate [51, 52, 38, 47]. Furthermore, exploiting the regularity of the Gaussian Process enables the agent to determine the next state to be explored without the requirement of known constraints [53, 54, 55]. An analytical model is then used to derive a policy to reach that state. However, this approach is limited to simple systems, as safely reaching the target state for high-dimensional systems remains challenging. Our approach also exploits the knowledge of the dynamics model and the constraints to construct a safe action space. Unlike existing approaches, ATACOM does not require an initial safe policy, a backup policy, or solving a constrained optimization problem. Instead, ATACOM constructs a safe action space where all sampled actions are guaranteed to be safe, making it compatible with any RL algorithm.
@@
Safe learning for robotics, as the core intersection between the SafeRL and robotics [56], has been raised for various types of application scenarios, varying from manipulation [57, 58] and navigation [59, 60], to locomotion [61] and Human-Robot Interaction (HRI) [62, 63, 64, 65]. Previous work of ATACOM has demonstrated the effectiveness of the method in manipulation, navigation, and HRI tasks [12, 13]. In this paper, we conduct a rigorous theoretical analysis, provide clear illustrations of individual components using simple low-dimensional environments, and ultimately validate the method through its application in a high-dimensional, real-world robotic task.
III Preliminaries and Problem Statement
Differential Geometry
@@ III Preliminaries and Problem Statement
$\mathrm{T}_{x}\mathcal{M}$ =$=\left\{{\bm{v}}\in\mathcal{E}:\langle\mathrm{grad}f_{i}({\bm{x}}),{\bm{v}}\rangle=0,i\in\{1,\dots,k\}\right\}$
The dimension of the tangent space is $\mathrm{dim}{\mathcal{M}}=d-k$.
LaSalle’s Invariance Principle
We present the LaSalle’s Invariance Principle [67], used later in the discussion to prove the safety. Consider the autonomous system
@@
Problem Statement
In SafeRL, we model the environment as a CMDP. A CMDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},P,\gamma,R,\mu_{0},\mathcal{K}\rangle$ with the state space $\mathcal{S}$, the action space $\mathcal{A}$, the state transition probability kernel $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{+}$, the discount factor $\gamma\in(0,1]$, the reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, the initial state distribution $\mu_{0}:\mathcal{S}\rightarrow\mathbb{R}^{+}$, and the set of constraint functions $\mathcal{K}:=\{k_{i}:\mathcal{S}\rightarrow\mathbb{R},i\in\{1,\dots,K\}\}$. In this paper, we approach the safety problem in RL from the point of view of Safe Exploration. The objective is to prevent constraint violations throughout the whole learning process, considering the following constrained optimization problem:
@@
with the trajectory $\tau=[{\bm{s}}_{0},{\bm{a}}_{0},\dots,{\bm{s}}_{T}]$, the state ${\bm{s}}_{t}\in\mathcal{S}\subset\mathbb{R}^{S}$, and the action ${\bm{a}}_{t}\in\mathcal{A}\subset\mathbb{R}^{A}$ sampled from a policy $\pi$. The inequality constraints $k_{i}:\mathcal{S}\rightarrow\mathbb{R}$ specify the safety requirements at each time step among all trajectories.
Modern RL algorithms solve this optimization by evaluating a sequence of policies $\{\pi_{0},\dots\pi_{j}\}$ in the environment, where the learning algorithm produces a new policy $\pi_{j+1}$ using $\mathcal{D}_{j}$, i.e. the dataset of environment interactions using the previous policies $\{\pi_{0},\dots,\pi_{j}\}$. In this scenario, Safe Exploration requires that every state ${\bm{s}}$ from every trajectory $\tau$ sampled from each policy $\pi_{j}$ is safe, i.e., $k_{i}({\bm{s}})\leq 0,\forall i$. This requires that the optimization algorithm selects each policy $\pi_{j}$ from $\Pi_{\text{safe}}$, i.e., the space of policies respecting the safety constraints. Most safe exploration approaches design a space $\tilde{\Pi}\subseteq\Pi_{\text{safe}}$ and solve the optimization problem in (3) generating $\pi_{j}\in\tilde{\Pi}$.
In the rest of this work, we focus on the control affine system that is local Lipschitz continuous, described as follows:
@@ IV Methods
In this section, we introduce the technical details of our proposed approach Acting on the TAngent Space of the COnstraint Manifold (ATACOM) and discuss under which conditions the designed controller is safe. Briefly, our approach tries to construct a safe action space tangent to the Constraint Manifold. Combining the basis of the tangent space with a coordinate determined by the controller, that can be hand-crafted or an arbitrary RL method, we can guarantee the safety of the overall control action. The conceptual illustration of the approach is shown in Fig. 2. Our approach is particularly suitable for RL as the actions are sampled randomly from a safe space, and the policy is safe by construction.
As it is fundamental for our approach, we will first rigorously define the Constraint Manifold and the Tangent Space at each point on the manifold. Then, we introduce an augmented state dynamic model that incorporates the dynamics of the slack variable. Next, we derive a safe controller that moves on the constraint manifold equipped with augmented dynamics. We then theoretically analyze the conditions of the controller to guarantee safety. At the end of this section, we extend our safety proof from a Lipschitz continuous system to switched systems, as actions drawn from a stochastic policy in the RL settings break the assumption of the Lipschitz continuity.
@@ IV-A Definition of the Constraint Manifold
Constraint function $k$ is of class $C^{1}$.
@@
Consider the inequality constraint $k(s)\coloneqq s\leq 0$. The constraint manifold =$\mathcal{M}=\{(s,\mu)\in\mathbb{R}\times[0,\infty):c(s,\mu)\coloneqq s+\mu=0\}$ is a 1-dimensional submanifold embedded in $\mathbb{R}^{2}$. The constraint Jacobian is $J_{c}(s,\mu)=[1\quad 1]$. We can choose the tangent space basis as $B(s,\mu)=\frac{1}{\sqrt{2}}[1,-1]^{\intercal}$.